OpenStudy (pphalke):

John borrowed $1200 from his credit union at 18% simple interest. He agrees to repay the loan by making 30 equal monthly payments. 11. How much is the finance charge? (Remember interest rate and time must have matching units.) 12. How much is the monthly payment?

OpenStudy (dumbcow):

18% is yearly rate, convert to monthly by dividing by 12 ---> 18/12 = 1.5% Total Interest = 1200*30*.015 Monthly payment = (1200 + Total Interest)/30